In 5 years time, I will be 3 times as old as my son. In 14 years time, the sum of our ages will be 62. Today the sum of our ages is?
Answer by jorel1380(3719)

m+5=3(s+5)
m+14+s+14=62
----------------
m+s+28=62
m+s=34
m=3(s+5)-5
m=34-s
3(s+5)-5=34-s
3s+15-5+s=34
4s+10=34
4s=24
s=6
m=28
I'm 28. My son is 6. The sum of our ages is 34..